A security code consist of two letter followed by two digits. The first letter cannot be A, B or C. What is the probability of guessing the security code if you are allowed 500 tries.
Answer #1

A security code consist of two letter followed by two digits

letters = 26

digits = 10

Fundamental counting,

first letter cannot be A, B or C = 23

Second letter = 26

Third digit = 10

Fourth digit = 10

23*26*10*10 = 59800

the probability of guessing the security code if you are allowed 500 tries = 500/59800 = 0.0084